Find g(x,x),g(x,y),g(y,x) for the following function g(x,y)=ax^2+bxy+cyz^2

Answer by Edwin McCravy(20054)

g(x,y)=ax^2+bxy+cyz^2

To find g(x,x), substitute x for y and simplify

g(x,x)=ax^2+bxx+cxz^2

g(x,x)=ax^2+bx^2+cxz^2

To find g(x,y), copy it over

g(x,y)=ax^2+bxy+cyz^2

To find g(y,x), substitute x for y and y for x:

g(y,x)=ay^2+byx+cyz^2

Edwin